OpenStudy (ffasinger):

What is the equation of the following line? Be sure to scroll down first to see all answer options.

OpenStudy (ffasinger):

A. y = -3x B. y = x C. y = 3x D. y = - x E. y = -4x F. y = - x

OpenStudy (ffasinger):

b is y=1/4x Dy=- 3/4x F y=3/4x

OpenStudy (anonymous):

from \((0,0)\) to \((4,-3)\) how many units to the right from \(0\) to \(4\)?

OpenStudy (anonymous):

it is kind of a silly question since the answer is already given to you, but that is what you need

OpenStudy (ffasinger):

so what would it be i am even more confused???

OpenStudy (anonymous):

from 0 to 4 is how far?

OpenStudy (anonymous):

actually forget that , since the line goes through \((0,0)\) the number in front of the x (the coefficient) is the second coordinate over the first coordinate of the point \((4,-3)\)

OpenStudy (ffasinger):

so it would be F i think

OpenStudy (anonymous):

i can't read those it is the one with the coefficient in front of the x that is \(\frac{-3}{4}\)

OpenStudy (anonymous):

oh i see, no, not F, D keep the minus sign

OpenStudy (ffasinger):

yeah that what it is

OpenStudy (ffasinger):

@satellite73

OpenStudy (ffasinger):

@That.Asian.Kid.

OpenStudy (anonymous):

yes, D